Conner is camping along a river. It takes him 1.5 hours to paddle his canoe 6 miles upstream from his campsite. Conner turns his canoe around and returns 6 miles downstream to his campsite in exactly 1 hour. What is the rate of the river's current and the rate of conner's paddling in still water?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

boat speed =x mph
current speed =y mph
against current x-y 1.50 hours
with current x+y 1.00 hours

Distance = same= 6 miles
t=d/r
6/(x-y)= 1.50
1.5(x-y)= 6.00
1.5x-1.5y= 6 ....................1
6/(x+y )= 1.00
1.00(x+ y)=6
1.00x+1.00y=6 ...............2
Multiply (1) by 1.00
Multiply (2) by 1.50
we get
1.5x-1.5y=6
1.5 x + 1.5 y = 9
3 x = 15
/ 3
x = 5 mph

Answer by stanbon(75887)

Conner is camping along a river. It takes him 1.5 hours to paddle his canoe 6 miles upstream from his campsite. Conner turns his canoe around and returns 6 miles downstream to his campsite in exactly 1 hour. What is the rate of the river's current and the rate of conner's paddling in still water?
-----------------------
Upstream DATA:
time = 1.5 hrs ; distance = 6 miles ; rate = d/t = 6/1.5 = 4 mph
-----
Downstream DATA:
distance = 6 miles ; time = 1 mr ; rate = d/t = 6/1 = 6 mph
-----
Equations:
Upstream::: b - c = 4 mph
Downstream b + c = 6 mph
---
Add and solve for "b":
2b = 10 mph
b = 5 mph (speed of the boat in still water)
---
Solve for "c":
b + c = 6
5 + c = 6
c = 1 mph (speed of the current)
====================================
Cheers,
stan H.